Can you help me with this problem?
Solve by elimination: -5y+2x=4
-0.02y+0.03x=0.04
Thanks

Answer by Alan3354(69443)

-5y+2x=4 --> -15y+6x=12 (x3)
-0.02y+0.03x=0.04 --> -4y+6x=8 (x200)
-15y+6x=12
-4y +6x=8 Subtract
----------
-11y = 4
y = -4/11 yyyyyyyyyyyyyyyyy
----------
-5y+2x=4
-5(-4/11) + 2x = 4
20/11 + 2x = 4
2x = 24/11
x = 12/11 xxxxxxxxxxxxxxxx